# Towards Interpretable Deep Generative Models via Causal Representation Learning

### v1

@@ 3.1 Identifiability assumptions on the loadings matrix
Triangular loadings matrices.  One early approach to identify the linear factor analysis model was to constrain the class of loadings matrices to be lower triangular matrices with positive diagonal elements (PLT) (Anderson and Rubin,, 1956). There are no rotation matrices that preserve this PLT structure except the identity, so $\bm{B}$ is identifiable. The PLT assumption is commonly used in the statistical factor analysis literature (Carvalho et al.,, 2008; Frühwirth-Schnatter and Lopes,, 2010).
While the PLT identifies $\bm{B}$, it does so at the expense of “fixing” the factors to the first $K$ dimensions of $\bm{X}$. That is, with the triangular assumption, the generative model is:
@@
Sparsity and anchor features.
More recently, Bing et al., (2020) proposed a method which learns which observed features should be used as “anchors” for the factor dimensions. Bing et al., (2020) prove their method identifies the loadings matrix with the key assumption of “pure variables”, or what we call “anchor features”. An anchor feature for a particular factor is an observed feature that depends only on this factor. In other words, no other factor influences this feature. The anchor features assumption of Bing et al., (2020) then demands that for every factor dimension, there are at least two anchor features. The anchor features do not need to be known in advance; these features have observable implications in the data covariance and thus can be inferred (see Figure 2(b) for a visualization).
The anchor features assumption means that the true loadings matrix $\bm{B}$ contains two submatrices that are the identity (after row and column permutation):
@@
The anchor feature assumption has also been utilized for identification in related latent variable models. For example, this is analogous to the “anchor word” assumption in topic models (Arora et al.,, 2013). This is also related to the separability condition (Donoho and Stodden,, 2003) in the non-negative matrix factorization (NMF) literature.
The anchor feature and separability assumptions also help explain the empirical success of sparse Bayesian factor analysis. In Bayesian factor analysis, sparsity-inducing priors are often placed on the loadings matrix $\bm{B}$; these sparse models often yield interpretable groups of features in real data settings (Knowles and Ghahramani,, 2011; Bhattacharya and Dunson,, 2011; Ročková and George,, 2016; Wang and Stephens,, 2021). These sparsity-inducing priors will place higher probability on loadings matrices that satisfy the anchor feature assumption assumption than dense loadings matrices.
@@ 3.2 Identifiability assumptions on the distribution of the factors
Heavy-tailed factors.  Rohe and Zeng, (2023) show that the factors are identifiable if their distribution belongs to a class of heavy-tailed distributions. These heavy-tailed distributions are not rotationally invariant, thus limiting the set of equivalent-likelihood solutions. [^1] Maxwell’s theorem states that the multivariate Gaussian is the only rotationally invariant distribution of independent variables (Maxwell,, 1860 and III, 4 in Feller,, 1971.) The heavy-tailed assumption on the factors also has precedent in the literature on independent component analysis (ICA, Comon,, 1994). Another key aspect of Rohe and Zeng, (2023) is that they provide theoretical justification for a long-used technique: principal component analysis (PCA) followed by a Varimax rotation (Kaiser,, 1958). Specifically, Rohe and Zeng, (2023) prove this “vintage” procedure consistently estimates the latent factors.
Correlated factors.
Typically the latent factors are assumed to be uncorrelated, but recent work has begun to investigate the case where factors are correlated: Bing et al., (2020) allow for correlated factors, as long as the variance of each factor is greater than its covariance with any other factor; Rohe and Zeng, (2023) implicitly allow for correlated factors by estimating a “middle matrix” between the loading and the factor matrices.
@@ 3.3 Beyond classical factor analysis
  • The PLT assumption is equivalent to assuming a known causal ordering on the factors;
  • Anchor features are equivalent to assuming a particular sparsity pattern on the graphical model implied by the loadings matrix $\bm{B}$;
  • The use of heavy-tailed factors is closely related to non-Gaussianity assumptions that have been used to identify ICA models;
  • Correlated factors arise naturally with causally related latent factors, and assumptions on correlated factors are closely related to causal restrictions on the latent factors.
Silva et al., (2006) recognized this connection with causal models, and presented a framework for learning causal structure in linear latent variable models. For example, can we determine which biological processes are causally related? In essence, this is the goal of causal representation learning: How do we learn latent factors (i.e. representations), and learn the causal relationships between them? To study such questions, we need some concepts some the graphical approach to causality (Pearl,, 2000), which we now review.
@@ 4 Causal Models and Latent Variables
Evidently, there is a close connection between classical identifiability assumptions in factor analysis and graphical assumptions that arise in machine learning and causal applications. In fact, the usual factor model is equivalent to a particular type of graphical model, sometimes called a measurement model (Silva et al.,, 2006), where there are no edges between observed variables, or from observed variables to latent variables. This relates to the factor analysis model where the $\bm{x}_{i}$ are independent conditional on $\bm{z}_{i}$.
In order to keep this review self-contained, in this section we review some minimal background on graphical models, with a focus on the linear setting to allow for easy comparison with the previous section. Readers already familiar with graphical models will immediately recognize that everything extends to nonlinear models and in subsequent sections, we will consider general nonlinear models.
@@ 4.1 Graphical approach to causality
The graphical approach to causality is built upon the fundamental concept of a Bayesian network. A Bayesian network (BN) over a random vector $\bm{x}_{i}=(\bm{x}_{i1},\ldots,\bm{x}_{iD})\in\mathbb{R}^{D}$ with joint distribution $P$ is any directed acyclic graph (DAG) $\mathsf{B}=(\mathsf{V},\mathsf{E})$ with $\mathsf{V}=\bm{x}_{i}$ such that $P$ factorizes over $\mathsf{B}$, i.e. it satisfies the Markov condition or is Markov to $\mathsf{B}$:
@@
[…]
We are interested in models that contain both observed and hidden components: $(\bm{x}_{i},\bm{z}_{i})=(\bm{x}_{i1},\dots,\bm{x}_{iD},\bm{z}_{i1},\dots,\bm{z}_{iK})$. In this case, the DAG $\mathsf{B}$ and the factorization (4.1) extends over the joint random vector $(\bm{x}_{i},\bm{z}_{i})$. Without some type of structural assumption on $\mathsf{B}$, the types of structures that can be represented by a Bayesian network with latent variables are known to be quite exotic, and have been the subject of much research (Evans,, 2016; Evans et al.,, 2018; Richardson et al.,, 2017). We will focus our review on the case of factor structures, defined below, in order to emphasize the connection with factor analysis, although one could of course study more exotic latent structures.
Factor models as in (3.1) implicitly impose structural assumptions in terms of the DAG $\mathsf{B}$ (Drton et al.,, 2007, 2023). That is, if $\mathsf{B}$ is a BN for a factor model over $(\bm{x}_{i},\bm{z}_{i})$, then $\mathsf{B}$ has a factor structure defined as follows:
@@
This definition amounts to the assumption that $\bm{x}_{i}$ are conditionally independent given $\bm{z}_{i}$, as in the factor analysis model. In fact, if we assume linear relationships, Definition 1 is equivalent to (3.1). Figure 1 depicts two prototypical examples of a factor model.
@@ 4.2 Latent causal graphs
More formally, factor models always decompose into two parts (Figure 4): A bipartite component that involves only directed edges from $\bm{z}_{i}\to\bm{x}_{i}$, and a latent causal graph $\mathsf{G}$ that involves only directed edges between the latent variables $\bm{z}_{i}$. Thus, the full causal graph over $(\bm{x}_{i},\bm{z}_{i})$ jointly is denoted by $\mathsf{B}$, whereas the latent causal graph over $\bm{z}_{i}$ only is denoted by $\mathsf{G}$. We omit reference to the bipartite graph since this is not the main focus. See Figure 4.
$z_{1}$$z_{2}$$z_{3}$$z_{4}$222233-3- 333331111

### v2

@@ 3.1 Identifiability assumptions on the loadings matrix
Triangular loadings matrices.  One early approach to identify the linear factor analysis model was to constrain the class of loadings matrices to be lower triangular matrices with positive diagonal elements (PLT) (Anderson and Rubin, 1956). There are no rotation matrices that preserve this PLT structure except the identity, which limits the set of equivalent-likelihood solutions. The PLT condition is commonly used in the statistical factor analysis literature (Carvalho et al., 2008; Frühwirth-Schnatter and Lopes, 2010).
An issue with the PLT condition, however, is that it “fixes” the factors to the first $K$ dimensions of $\bm{X}$. That is, with the triangular assumption, the generative model is:
@@
More recently, Frühwirth-Schnatter et al. (2023) generalized the PLT condition to the unordered generalized lower triangular (UGLT) condition, proving that the UGLT condition resolves unidentifiability due to rotational invariance. The UGLT condition allows for an unknown set of $K$ (out of $D$) features to have special status in defining the factors (instead of the first $K$ features, as in the PLT).
Anchor features.  Bing et al. (2020) propose learning which observed features act as “anchors” for the factor dimensions. Their method identifies the loadings matrix under the key assumption that each factor has at least two “pure variables”, or what we call “anchor features”. An anchor feature for a particular factor is an observed feature that depends only on that factor. Anchor features do not need to be known in advance; they can be inferred from the data covariance, under additional assumptions. See Figure˜2(b) for a visualization—anchor features exhibit higher mutual covariance than non-anchor features.
The anchor feature assumption means that the true loadings matrix $\bm{B}$ contains two submatrices that are the identity (after row and column permutation):
@@
The anchor feature assumption has also been utilized for identification in related latent variable models. For example, this is analogous to the “anchor word” assumption in topic models (Arora et al., 2013). This is also related to the separability condition (Donoho and Stodden, 2003) in the non-negative matrix factorization (NMF) literature.
The anchor feature and separability assumptions also help explain the empirical success of sparse Bayesian factor analysis. In Bayesian factor analysis, sparsity-inducing priors are often placed on the loadings matrix $\bm{B}$; these sparse models often yield interpretable groups of features in real data settings (Knowles and Ghahramani, 2011; Bhattacharya and Dunson, 2011; Ročková and George, 2016; Wang and Stephens, 2021). Such priors place higher probability on sparse matrices, which are more likely to satisfy the anchor assumption than dense matrices.
@@ 3.2 Identifiability assumptions on the distribution of the factors
Non-Gaussian independent factors.  In psychometrics and statistics, a popular approach to factor analysis is to first calculate principal components or maximum likelihood estimates, and then apply a rotation such as Varimax (Kaiser, 1958). (See for example, multivariate analysis textbooks such as Anderson (2003); Jolliffe (2002).) However, this approach lacked theoretical justification until recently, when Rohe and Zeng (2023) proved that PCA and Varimax consistently estimates the latent factors (provided the factors have a particular heavy-tailed distribution). Intuitively, Varimax rotates factors toward non-Gaussian solutions. As non-Gaussian distributions of independent variables are not rotationally invariant (as known from Maxwell’s theorem, Maxwell, 1860; see also III, 4 in Feller, 1971), this assumption removes indeterminances due to rotational invariance.
The relationship between non-Gaussianity and identifiability in factor models is also widely exploited in the literature on independent component analysis (ICA, Comon, 1994; Eriksson and Koivunen, 2004). Central to this literature is the Darmois-Skitovich lemma and its extensions, which extends Maxwell’s characterization of the normal distribution in terms of independent linear forms. See Kagan et al. (1973) for a detailed account and related results. An application to factor analysis is the following widely known identifiability result for ICA: If $\bm{x}_{i}=\bm{B}\bm{z}_{i}$, where $\bm{z}_{i}$ consists of independent non-Gaussian components, then $\bm{B}$ is identifiable up to permutation and scaling. This result crucially relies on non-Gaussianity, as even two Gaussian components is enough to break identifiability.
Correlated factors.  Typically the latent factors are assumed to be uncorrelated, but recent work has begun to investigate the case where factors are correlated: Bing et al. (2020) allow for correlated factors, as long as the variance of each factor is greater than its covariance with any other factor; Rohe and Zeng (2023) implicitly allow for correlated factors by estimating a “middle matrix” between the loading and the factor matrices.
@@ 3.3 Beyond classical factor analysis
  • The PLT assumption is equivalent to assuming a known causal ordering on the factors (e.g. Germain et al., 2015; Shen et al., 2022);
  • Anchor features are equivalent to assuming a particular sparsity pattern on the graphical model implied by the loadings matrix $\bm{B}$ (e.g. Silva et al., 2006; Arora et al., 2013);
  • Correlated factors arise naturally with causally related latent factors, and assumptions on correlated factors are closely related to causal restrictions on the latent factors (e.g. Silva et al., 2006).
Silva et al. (2006) made the connection between causality and factor models explicit by presenting a framework for learning causal structure in linear latent variable models. For example, can we determine which biological processes are causally related? In essence, this is the goal of CRL: How do we learn latent factors (i.e. representations), and learn the causal relationships between them? To study such questions, we need concepts from the graphical approach to causality (Pearl, 2009), which we now review.
@@ 4 Causal Models and Latent Variables
Evidently, there is a close connection between classical assumptions in factor analysis and graphical assumptions found in machine learning and causal inference. In fact, the factor analysis model is equivalent to a graphical model known as a measurement model, where the observed variables $\bm{x}_{i}$ are conditionally independent given latent variables $\bm{z}_{i}$. Measurement models originate from the psychometric literature on confirmatory factor analysis, where the causal relationships among latent variables are typically specified beforehand (e.g. Bentler, 1980; Bollen, 1989). By contrast, the contemporary measurement model used in CRL—where both the latent variables and their causal structure are unknown and must be inferred from data—was first explicitly formulated by Silva et al. (2006). Although this early work focused on linear models, modern advances increasingly emphasize nonlinear models.
For context, we provide some formal background on these notions in this section. In doing so, we will also present some minimal background on graphical models, emphasizing linear models for straightforward comparison with the factor analytic frameworks discussed above. However, readers familiar with graphical models will recognize that these concepts immediately generalize to nonlinear models.
@@ 4.1 Graphical approach to causality
The graphical approach to causality is built upon the fundamental concept of a Bayesian network. A Bayesian network (BN) over a random vector $\bm{x}_{i}=(x_{i1},\ldots,x_{iD})\in\mathbb{R}^{D}$ with joint distribution $P$ is any directed acyclic graph (DAG) $\mathsf{B}=(\mathsf{V},\mathsf{E})$ with $\mathsf{V}=\bm{x}_{i}$ such that $P$ factorizes over $\mathsf{B}$, i.e. it satisfies the Markov condition or is Markov to $\mathsf{B}$:
@@
The dependence on $\mathsf{B}$ in the above expression arises from the parent sets $\operatorname{pa}_{\mathsf{B}}(j)\mathrel{\mathop{\ordinarycolon}}=\{k\mathrel{\mathop{\ordinarycolon}}\bm{x}_{ik}\to\bm{x}_{ij}\text{ in }\mathsf{B}\}$. As is conventional in graphical models, we will often simply write $k$ in place of $\bm{x}_{ik}$, e.g. $k\to j$ indicates an edge pointing from $\bm{x}_{ik}$ to $\bm{x}_{ij}$. The purpose of the Markov condition is to connect distributions and graphs: Although every distribution is Markov to some DAG, the Markov condition allows us to define statistical models through structural assumptions on distributions.
We are interested in models that contain both observed and hidden components: $(\bm{x}_{i},\bm{z}_{i})=(x_{i1},\dots,x_{iD},z_{i1},\dots,z_{iK})$. In this case, the DAG $\mathsf{B}$ and the factorization (4.1) extends over the joint random vector $(\bm{x}_{i},\bm{z}_{i})$. Without some type of structural assumption on $\mathsf{B}$, the types of structures that can be represented by a Bayesian network with latent variables are known to be quite exotic, and have been the subject of much research (Richardson et al., 2017; Evans and others, 2018). We will focus on the case of factor structures, defined below, in order to emphasize the connection with factor analysis, although one could of course study more exotic latent structures.
Factor models as in (3.1) implicitly impose structural assumptions in terms of the DAG $\mathsf{B}$ (Drton et al., 2007). That is, if $\mathsf{B}$ is a BN for a factor model over $(\bm{x}_{i},\bm{z}_{i})$, then $\mathsf{B}$ has a factor structure defined as follows:
@@
Factor models are also known as measurement models; for consistency of presentation, we stick to the term factor structure, acknowledging that this is synonymous with earlier usage of measurement model. This definition implies that $\bm{x}_{i}$ are conditionally independent given $\bm{z}_{i}$, as in a factor model. In fact, if we assume linear relationships and Gaussian noise, Definition 1 is equivalent to (3.1). Figure 1 depicts two prototypical examples of a factor model.
@@ 4.2 Latent causal graphs
More formally, factor models always decompose into two parts (Figure˜4): A bipartite component that involves only directed edges from $\bm{z}_{i}\to\bm{x}_{i}$, and a latent causal graph $\mathsf{G}$ that involves only directed edges between the latent variables $\bm{z}_{i}$. Thus, the full causal graph over $(\bm{x}_{i},\bm{z}_{i})$ jointly is denoted by $\mathsf{B}$, whereas the latent causal graph over $\bm{z}_{i}$ only is denoted by $\mathsf{G}$. We omit reference to the bipartite graph since this is not the main focus.
$z_{1}$$z_{2}$$z_{3}$$z_{4}$223-33311
$\left.\begin{aligned} z_{i1}&=\nu_{i1}\\ z_{i2}&=2z_{i1}+\nu_{i2}\\ z_{i3}&=-3z_{i1}+3z_{i2}+\nu_{i3}\\ z_{i4}&=z_{i3}+\nu_{i4}\end{aligned}\right\}\Rightarrow\bm{z}_{i}=\bm{A}\bm{z}_{i}+\bm{\nu}_{i},\quad\bm{A}=\begin{pmatrix}0&0&0&0\\ 2&0&0&0\\ -3&3&0&0\\ 0&0&1&0\\ \end{pmatrix}$
In the linear case, also known as a linear structural equation model (SEM, Bollen, 1989; Pearl, 2009), the model for $p_{\bm{z}}(\bm{z}_{i})$ becomes
$\bm{z}_{i}=\bm{A}\bm{z}_{i}+\bm{\nu}_{i},\quad\bm{A}=(a_{jk})\in\mathbb{R}^{K\times K},$ (4.2)
where $\bm{\nu}_{i}$ are jointly independent, zero-mean random variables. Then we can associate to $\bm{z}_{i}$ a latent causal graph $\mathsf{G}$ by reading off the nonzero entries of the matrix $\bm{A}$: Simply define $\mathsf{G}=(\mathsf{V},\mathsf{E})$ by
$\mathsf{V}=\bm{z}_{i},\quad\mathsf{E}=\{(j,k)\mathrel{\mathop{\ordinarycolon}}a_{jk}\neq 0\}.$ (4.3)
See Figure˜5 for an example. It turns out that as long as $\mathsf{G}$ is acyclic, then $\mathsf{G}$ is a BN for $\bm{z}_{i}$. In other words, the matrix $\bm{A}$ explicitly encodes a BN for $\bm{z}_{i}$.
4.3 Interventions and causality
The edges in a DAG can be interpreted as indicating the causal dependencies in a system: The edge $z_{ik}\to z_{ij}$ indicates that $z_{ij}$ depends causally on $z_{ik}$ in the sense that changes to $z_{ik}$ propagate to $z_{ij}$. Returning to the linear SEM example in (4.2), such causal dependencies can be seen by writing the model component-wise:
$z_{ij}=\sum_{k=1}^{K}a_{jk}z_{ik}+\nu_{ij}.$ (4.4)
Explicitly, $z_{ij}$ is a noisy function of its parents.
Interventions and independent mechanisms.  This causal interpretation allows us to reason about how the system changes when we externally manipulate the values of certain variables, i.e. an intervention. An intervention occurs when we fix the value of a certain variable (or set of variables) to a particular value. When this happens, the causal parents of the intervened variable are ignored, akin to deleting incoming edges to this node.
To make this concrete, consider the linear SEM in Figure˜5. Without any external interventions, the value of $z_{i3}$ is determined by the values of $z_{i1}$ and $z_{i2}$. The value of $z_{i2}$, in turn, is determined by $z_{i1}$, whereas $z_{i1}$ is an independent exogenous variable that is not determined by any other $z_{ik}$. Now suppose we intervene on $z_{i2}$: Then the edge from $z_{i1}\to z_{i2}$ is broken, that is, the value of $z_{i1}$ is no longer relevant for determining $z_{i2}$: The value of $z_{i2}$ is fixed by external manipulation. This then propagates down to $z_{i3}$, whose value is determined by the (unintervened) value of $z_{i1}$ and the intervened value of $\bm{z}_{i2}$. The result is a new linear SEM
$z_{i1}$ $=\nu_{i1}$
$z_{i2}$ $=z_{i2}^{*}$
$z_{i3}$ $=-3z_{i1}+3z_{i2}^{*}+\nu_{i3}$
$z_{i4}$ $=z_{i3}+\nu_{i4},$
where $z_{i2}^{*}$ is the value that $z_{i2}$ is manipulated to. To denote the new distribution of $z_{i3}$ given this intervention on $z_{i2}$, the convention is to use do-notation:
$P(z_{i3}|z_{i1},\mathrm{do}(z_{i2}=z_{2}^{*})).$
Implicit in this formalization of interventions is the principle of modularity or independent mechanisms: Intervening on a node changes the local conditional distribution of that node, but leaves the remaining conditional distributions unchanged. See Peters et al. (2017) (Chapter 2) for more discussion on the principle of independent mechanisms.
Latent interventions.  The factor model (Definition˜1) makes explicit that intervening on the observed features $\bm{x}_{i}$ is not meaningful. Instead, we are interested in intervening on latent variables, which might seem strange at first glance. After all, we do not directly observe latent variables. In many applications, however, this is quite natural: recall the motivating examples in Section˜2.
When the observed features are pixels (Figure˜1(b)), intervening on individual pixels is not meaningful outside of toy applications. In practice, however, we can obtain additional images or data augmentations in which the object of the image is translated or rotated, or the background is changed. Each of these changes can be viewed as an intervention. However, it is difficult to model such interventions at the pixel level; e.g. in different images, different subsets of pixel values correspond to the background. A solution to this problem is to learn a latent representation $\bm{z}_{i}$ of each image $\bm{x}_{i}$ where the dimensions $z_{ik}$ have a consistent interpretation for all $i$; e.g. $z_{ik}$ corresponds to the rotation of the images. This type of data is readily available in applications, and can even be easily generated with the aid of simulators (e.g. Von Kügelgen et al., 2021; Brehmer et al., 2022).
In genomics, the observations are gene expression levels (Figure˜1(a)). Genes may coordinate as part of a biological pathway; collectively, these genes may induce effects in other variables. As such, it may be more efficient to aggregate these gene subsets into single causal variables representing gene programs or biological processes. Other examples of latent interventions include medical treatments (drugs, surgeries), radiation exposure, or other epigenetic factors (DNA methylation, histone modification, environmental factors).
We refer to an intervention where a single dimension of $\bm{z}_{i}$ is intervened upon as a single-node intervention. For a linear SEM, we formalize this as follows.
Definition 2 (Single-node latent interventions).
Given an intervention $e\in\mathcal{E}$, denote the resulting distribution on $\bm{z}$ as $p_{\bm{z}}^{(e)}$. An interventional distribution $p_{\bm{z}}^{(e)}$ is a single-node intervention of the latent distribution $p_{\bm{z}}$ if there is a single target node $t(e)\in[K]$, and the assigning equation of $z^{(e)}_{i,t(e)}$ is:
$z_{i,t(e)}^{(e)}$ $=\sum_{k=1}^{K}a^{(e)}_{t(e),k}z_{ik}^{(e)}+\nu_{i,t(e)}+c^{(e)}$ (4.5)
while leaving all other equations $z_{ij}$ for $j\neq t(e)$ unchanged. We assume $a_{t(e),k}^{(e)}\neq 0$ only if $k\to t(e)$ in $\mathsf{G}$; that is, no parents are added. The variable $c^{(e)}$ denotes a shift of the mean; the variance of the noise $\nu_{i,t(e)}$ may also be different to that of $p_{\bm{z}}$. An intervention is called perfect if $a_{t(e),k}^{(e)}=0$ for all $k\in[K]$, i.e. all connections to former parents are removed.
Intervention targets can be either known or unknown: When intervention targets are known, there is no ambiguity about the mapping $t(e)$ defined above. A considerably more challenging setting arises when intervention targets $t(e)$ are unknown. In CRL, since the intervention targets are latent and unobserved, these interventions are also unknown—in fact, we may not even know the correct number of latent factors to begin with.
This leads to additional complications in CRL: Although there is a growing literature on learning under unknown interventions (Eaton and Murphy, 2007; Kocaoglu et al., 2019; Jaber et al., 2020; Squires et al., 2020; Castelletti and Peluso, 2023; Hägele et al., 2022; Faria et al., 2022), this line of work focuses on interventions on measured variables (i.e. $x_{ij}$). When interventions are unknown and latent (i.e. $z_{ij}$), much less is known—this is a developing area of research driven by problems in CRL; e.g. Jiang and Aragam (2023).
5 What is causal representation learning?
With these preliminaries on factor models and causality out of the way, we are ready to discuss CRL in more precise terms. Simply put, CRL takes a modern view on factor analysis, driven by the success of generative AI and guided by three core principles as outlined in the introduction: Flexibility, interpretability, and transferability.
  • Flexibility is achieved by allowing (a) general nonlinear relationships through arbitrarily complex neural networks, as well as (b) correlations between latent variables;
  • Transferability is achieved by using causal models, which are widely believed to yield stable, invariant, transferable predictions across changing contexts and environments.
To capture these three core principles, CRL begins with a general nonlinear factor model of the form $\bm{x}_{i}=f(\bm{z}_{i})+\bm{\varepsilon}_{i}$, where the latent $\bm{z}_{i}$ are related by some latent causal graph $\mathsf{G}$ and the nonlinearity $f$ is typically estimated with deep neural networks, as in a deep generative model, and causal relationships are inferred through the use of data arising from multiple environments (e.g. across different experimental interventions, time points, etc.). Examples of generative models explicitly of this form include diffusion models, variational autoencoders, and flows. This class of models continues to dominate leaderboards tracking progress in generative AI (HuggingFace, 2025; Liu et al., 2024).
5.1 Problem definition
To discuss issues such as identifiability, we now formalize the CRL problem. Given a function class $\mathcal{F}$ and a family of prior distributions $\mathcal{P}$, we assume the observed data $\bm{x}_{i}\in\mathcal{X}$ is generated from the latent $\bm{z}_{i}\in\mathcal{Z}$ via an unknown mixing function $f\in\mathcal{F}$:
$\bm{x}_{i}=f(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\stackrel{{\scriptstyle\textup{iid}}}{{\sim}}p(\bm{\varepsilon}),\quad i=1,\dots,n,$ (5.1)
where the latent $\bm{z}_{i}\sim p_{\bm{z}}\in\mathcal{P}$ are generated according to a linear structural equation model:
$\bm{z}_{i}=\bm{A}\bm{z}_{i}+\bm{\nu}_{i},$ (5.2)
where $\bm{A}$ is the SEM matrix for some DAG $\mathsf{G}$ (cf. (4.2)) and $\bm{\nu}_{i}$ are jointly independent, zero-mean random variables. These assumptions are not necessary, but will help simplify the discussion here; see Remark 3 below.
The goal is to infer the unknown function $f$ and the prior on the latents $p_{\bm{z}}$ including $\bm{A}$, which encodes the causal graph $\mathsf{G}$ of $\bm{z}$. The parameters $(f,p_{\bm{z}},\bm{A})$ are important for sampling from the model and understanding the effect of interventions on the observed data. In CRL, the motivation for learning these parameters is interpretability (via sparse structure in $\bm{A}$ or $f$) and transferability to new domains (via the causal mechanisms encoded by $\bm{A}$). In practice, it may be enough to identify certain functionals of the parameters $(f,p_{\bm{z}},\bm{A})$, such as a subgraph or the behavior of $f$ on subspaces of interest.
These parameters are specified by the class of transformations $\mathcal{F}$ and the family of distributions $\mathcal{P}$ over $\bm{z}$. For example, in the linear factor analysis model $\mathcal{F}$ is a class of linear maps and $\mathcal{P}$ is typically the family of Gaussian distributions. The overall model can be formalized as the collection of distributions induced by these families. More general models can be obtained by expanding $\mathcal{F}$ to include nonlinear transformations or $\mathcal{P}$ to include non-Gaussian distributions. Since the only difference between (5.1) and the earlier linear factor model (3.1) is the replacement of the linear map $\bm{B}\bm{z}_{i}$ with the nonlinear map $f(\bm{z}_{i})$, the former is a special case of the latter. Moreover, since even linear factor models are not identifiable, it follows that (5.1) is also not identifiable. A key goal of CRL is therefore to determine conditions under which this model becomes identifiable and estimable.
Remark 2.
The model (5.1) is an assumption on the distribution that can be made independently of whether neural networks are used to fit the data or not; in particular it is not necessary to assume that $f$ itself is a neural network. A common choice for $\mathcal{F}$ is the family of invertible, differentiable functions (i.e. diffeomorphisms); see Kivva et al. (2022) for a detailed discussion of the implications of various functional restrictions. Of course, in practice we are interested in using neural networks to fit this model nonparametrically.
Remark 3.
We focus on models where the latents satisfy a linear structural equation model and the noise is additive, although these assumptions are not necessary. The literature has studied more general models (e.g. Jiang and Aragam, 2023; Varici et al., 2023, 2024; von Kügelgen et al., 2023); see the end of Section 6 for a brief overview.
Remark 4.
The model (5.1) is a nonlinear factor model, which has a long history in both statistics and ML; see Amemiya and Yalcin (2001) for a review. See also Zhou et al. (2024); Xu et al. (2023) for recent developments. Although the mathematical models share similarities, what differentiates CRL from prior work on nonlinear factor analysis is the focus on deep learning and causality that drives much of CRL. For example, a key distinction arises from the latent SEM (5.2) and the causal, interventionist semantics it provides.
5.2 Generation vs. understanding
As noted in the introduction, in applications of generative AI there is a crucial distinction between the problems of generation and understanding that parallels the two cultures of statistical modeling. Since the vast majority of the literature on generative AI focuses on generation only, we pause for a moment to outline in more detail these two goals:
  1. 1.